OpenStudy (anonymous):

A particle is undergoing simple harmonic motion. The equation of the particle's displacement is given by s(t) = 4cos (2t + 5), where t time in sec. Find the velocity of the particle at time t.

OpenStudy (anonymous):

So... CosU=U'Sin U

OpenStudy (anonymous):

take derivative

OpenStudy (anonymous):

-2Sin(2t+5) am I right so far:/

OpenStudy (anonymous):

The velocity is the derivative of the displacement:\[v(t) = s'(t) = 4(\cos(2t+5))' = -4\sin(2t + 5)(2t)' = -8\sin(2t+5).\]

OpenStudy (anonymous):

-4Sin(2t+5 )(2) -8 Sin(2t+5)

OpenStudy (anonymous):

Wow. I cant believe i didnt get it. You make it seem so easy :P